OpenStudy (anonymous):

What is the simplest form of the expression? (Gonna write it in comments) :P

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{6}\ sqrt{2}+\sqrt{6}\]

OpenStudy (anonymous):

Suppose to be first 2 over the second 2

OpenStudy (anonymous):

I know 6 reduces to 3 xD

OpenStudy (anonymous):

What do u mean by sqrt2 after sqrt 6, Just multiplied by sqrt6?

OpenStudy (anonymous):

sqrt2 - sqrt6 _____________ sqrt2 + sqrt6 (Divided)

OpenStudy (anonymous):

Ohhh, Hint, Time it by the conjugate and simplify ((sqrt2 +sqrt6)/(sqrt2+sqrt6))